OpenStudy (anonymous):

√28+√63+√175 show work please

OpenStudy (mathstudent55):

Can you factor 28 into two numbers, one of which is a perfect square?

OpenStudy (anonymous):

7*4?

OpenStudy (mathstudent55):

In this case, the perfect squares are the squares of the natural numbers, 1, 2, 3, 4, ..., so they are 1, 4, 9, 16, ...

OpenStudy (mathstudent55):

Great.

OpenStudy (mathstudent55):

Now can you do the same for 63?

OpenStudy (anonymous):

7*9

OpenStudy (mathstudent55):

Great again. Now do the same for 175.

OpenStudy (anonymous):

7*25 I THINK

OpenStudy (mathstudent55):

Correct again. This is what we have now: \(\sqrt{28} + \sqrt{63} + \sqrt{175} \) \(=\sqrt{4 \times 7} + \sqrt{9 \times 7} + \sqrt{25 \times 7} \)

OpenStudy (anonymous):

what do I do after that

OpenStudy (pinksapphire):

Split them up first: \[(\sqrt{4}+\sqrt{7})+(\sqrt{7}+\sqrt{9})+(\sqrt{7}+\sqrt{25})\] \[2\sqrt{7}+3\sqrt{7}+5\sqrt{7}\] Add them up: \[10\sqrt{7}\]

OpenStudy (mathstudent55):

Now we separate the roots as follows: \(=\sqrt{4}\sqrt{7} + \sqrt{9} \sqrt {7} + \sqrt{25} \sqrt{ 7} \) Now take the square roots of all the perfect squares.

OpenStudy (mathstudent55):

@PinkSapphire You made a big mistake.

OpenStudy (pinksapphire):

What?

OpenStudy (mathstudent55):

\(=2\sqrt{7} + 3 \sqrt {7} + 5 \sqrt{ 7} \) Now treat all terms as like terms, since they all have a factor of \(\sqrt{7} \) and add them up.

OpenStudy (mathstudent55):

Compare your first line to my first line after yours.

OpenStudy (pinksapphire):

It's basically the same...

OpenStudy (anonymous):

where did the 2 befroe the √ come from?

OpenStudy (mathstudent55):

No. Your idea is good, but your mistake is that the perfect squares you split up are MULTIPLIED by the remaining parts, not added.

OpenStudy (pinksapphire):

No they are added...

OpenStudy (mathstudent55):

I did say take the square roots of the perfect squares. \(\sqrt{4} = 2\)

OpenStudy (anonymous):

but that only equals 15.8745

OpenStudy (pinksapphire):

Where did you get that from?

OpenStudy (anonymous):

as I just typed the original equation in my calculator I got 26.4575

OpenStudy (pinksapphire):

You can't do that...

OpenStudy (anonymous):

I know thats why I'm asking for help

OpenStudy (mathstudent55):

@Chrisgoblin Ignore this response. It's not correct. @PinkSapphire Read below: You wrote this (which is incorrect) \( (\sqrt{4}+\sqrt{7})+(\sqrt{7}+\sqrt{9})+(\sqrt{7}+\sqrt{25})\) If you continue correctly from it you'd get this: \( =(2+\sqrt{7})+(\sqrt{7}+3)+(\sqrt{7}+5)\) \( =2+ 3 + 5 + \sqrt{7}+\sqrt{7}+\sqrt{7}\) \(= 10 + 3\sqrt 7\) Which is not the same as the correct answer \(10 \sqrt{7} \)

OpenStudy (mathstudent55):

OK, @Chrisgoblin Let's continue.

OpenStudy (mathstudent55):

@Chrisgoblin Did you understand until this step? \(=\sqrt{4}\sqrt{7} + \sqrt{9} \sqrt {7} + \sqrt{25} \sqrt{ 7}\)

OpenStudy (anonymous):

yea

OpenStudy (mathstudent55):

Here's how we continue. We take the square root of each perfect square. I color coded it to make it easier to see how you go form one step to the next. \(=\color{red}{\sqrt{4}}\sqrt{7} + \color{green}{\sqrt{9}} \sqrt {7} + \color{blue}{\sqrt{25}} \sqrt{ 7}\) \(=\color{red}{2}\sqrt{7} + \color{green}{3} \sqrt {7} + \color{blue}{5} \sqrt{ 7}\)

OpenStudy (anonymous):

oh I see

OpenStudy (mathstudent55):

Now the last step is to add all the square roots of 7 together. Think of square root of 7 as x Since 2x + 3x + 5x = 10x, then similarly, \(2\sqrt{7} + 3 \sqrt {7} + 5 \sqrt{ 7} = 10\sqrt{7}\)

OpenStudy (anonymous):

26.4575

OpenStudy (anonymous):

x=√7 right?

OpenStudy (mathstudent55):

The answer \(10\sqrt 7\) is correct. If your problem is asking for an approximation, then you can use 26.4575, but my answer is exact. Also, if an approximation were what the problem wants, you could just take the square root of each number in the beginning and added them all up. There isn't much work to show then.

OpenStudy (mathstudent55):

Yes, \(x = \sqrt 7\)

OpenStudy (anonymous):

thanks and do you think you could help me with one more problem @mathstudent55

OpenStudy (mathstudent55):

You're welcome, and yes.

OpenStudy (anonymous):

(√2w+5)+4=√25

OpenStudy (mathstudent55):

The first one or the second one? (The difference is whetehr or not the w is also in the square root.) \((\sqrt{2}w + 5) + 4 = \sqrt{25} \) \((\sqrt{2w} + 5) + 4 = \sqrt{25} \)

OpenStudy (anonymous):

its the square root of 2w +5 the top line covers 2w+5

OpenStudy (anonymous):

thats why I put the parenthesis

OpenStudy (mathstudent55):

Oh, ok. Let me write that, then.

OpenStudy (mathstudent55):

You meant the parentheses after the root symbol. √(2w+5)+4=√25 \(\sqrt{2w + 5} + 4 = \sqrt{25} \) This is it now?

OpenStudy (anonymous):

oh yea woops thats it

OpenStudy (mathstudent55):

Ok. Let's solve it then.

OpenStudy (mathstudent55):

First, take the square root of 25 on the right side.

OpenStudy (mathstudent55):

\(\sqrt{2w + 5} + 4 = \color{red}{\sqrt{25}}\) \(\sqrt{2w + 5} + 4 = \color{red}{5}\)

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

Now subtract 4 from both sides. What do you get?

OpenStudy (anonymous):

√(2w+5)=1

OpenStudy (mathstudent55):

Good. Now square both sides.

OpenStudy (mathstudent55):

When you square a square root, just remove the square root symbol.

OpenStudy (anonymous):

ok 2w+5=1

OpenStudy (mathstudent55):

Great. Now just solve for w. Subtract 5 from both sides. Then divide both sides by 2.

OpenStudy (anonymous):

2w=-4

OpenStudy (mathstudent55):

Now divide both sides by 2.

OpenStudy (anonymous):

1w-2

OpenStudy (mathstudent55):

w = -2

OpenStudy (anonymous):

wow I feel stupid

OpenStudy (anonymous):

*w=-2 so √(2*-2+5)+4=√25

OpenStudy (anonymous):

so 5

OpenStudy (mathstudent55):

There is one more check. Sometimes when both sides of an equation are squared, extra solutions are introduced that are not solutions to the original equations. These extra solutions are called "extraneous solutions." We need to check our solution in our original equation. \(\sqrt{2w + 5} + 4 = \sqrt{25}\) \(\sqrt{2(-2) + 5} + 4 = 5\) \(\sqrt{-4 + 5} + 4 = 5\) \(\sqrt{1} + 4 = 5\) \(1 + 4 = 5\) \(5 = 5\) The solution is correct, w = -2.

OpenStudy (anonymous):

thanks